# GraspLDM: Generative 6-DoF Grasp Synthesis using Latent Diffusion Models

### v1

@@ IV Results and Discussion
We evaluate the performance and scaling of our GraspLDM on two category sets- 1C and 63C, where C denotes the number of ShapeNetSem categories from the ACRONYM dataset. 1C is composed of 110 Mugs in the train set and 50 held-out Mugs in the test set. The 63C set contains 1100 objects from 63 categories in the train set and 400 held-out objects in the test set. For observations, we use full object point clouds and partial point clouds. Full object point clouds are sampled from the mesh surface. On the other hand, partial point clouds are rendered from a virtual depth camera at a distance between 30cm and 1m. In both cases, point clouds are randomly rotated, noised, and regularized to 1024 points which is the input to our point cloud encoder.
@@
We evaluate success rate of the grasp poses generated by the models in a large-scale parallel simulation in Isaac Gym [21]. In the simulation, we use only a Franka-Emika two fingered gripper without the robotic arm and ignore gravity to avoid the effect of external factors on the evaluations. For each grasp pose, the simulation executes three steps to report a grasp success or a failure. (1) The gripper is spawned at the given grasp pose relative to the object with the fingers open. (2) The fingers are closed at the grasp pose. (3) The gripper lifts the object 1m in the +Z direction. Finally, a grasp is reported successful if the object remains attached to the gripper. Note that the evaluation is conservative as any penetration of the gripper in the first step or adverse contact in the second stage will result in a grasp failure. On the other hand, we observe that for many large objects, ground truth grasp poses in the ACRONYM dataset provide a very low or zero success rate. Therefore, we filter out the objects in each set for which the ground truth grasps fail more than 75% of the time.
@@
While the success rate evaluates the quality of grasp generation, a model with a high success rate may generate grasps only around a small region of the object, instead of covering all the graspable regions around the object. Consequently, we also use SE(3) Earth Mover’s Distance (EMD) metric [9] to evaluate how well the trained models learn the object-centric grasp pose distribution. The SE(3) EMD metric evaluates the empirical distance between two distributions of SE(3) poses. To compute this, we sample 100 grasps per object from a model and sample 100 random grasps from the ground truth datasets. To further validate the flexibility of GraspLDM, we provide additional results on sampling speed and task-conditional generation.
Figure 4: Visualization of latent space de-noising in GraspLDM. Reverse diffusion de-noises these latents from $t=T$ to $t=0$, gradually moving the bad grasps towards regions of good grasps.
@@
For success rate, we outline two observations in Fig. 2(a). First, the LDM models consistently improve upon the base VAE models, demonstrating that the sample quality improves by using a diffusion model in VAE’s latent space. Second, the success rate performance of GraspLDM models scales better to a larger object set (63C). Comparing the median of the success rate on the test set in Fig. 2(a), GraspLDM-1C improves the median success rate to 88% compared to 84% of GraspVAE-1C. On the larger 63C set, the GraspLDM-63C model improves the median success rate to 78% compared to 66% from GraspVAE-63C. GraspLDM models also improve the interquartile range in each case. For the comparison against the baseline, SE3-DiF-1C model reports a median success rate of 89% which is comparable to 88% the GraspLDM-1C model. On the larger 63C set, we observe that the GraspLDM-63C model registers 78% median success rate and distinctly higher interquartile range compared to SE3-DiF-63C. Therefore, GraspLDM models scale more favorably to larger object sets compared to the state-of-the-art approaches. During the tests, we also observed that the models available from [9] do not hold out a test split and is trained on all objects. Therefore, the comparison to SE3-DiF models [9] is conservative, which emphasizes the performance of GraspLDM on withheld objects.
@@
Overall, the results validate our hypothesis that a diffusion model in the latent space can bridge the prior gap in the latent space to provide higher-quality grasp pose samples. Fig. 4 visualizes this effect where the latent de-noising moves latents corresponding to bad grasps towards good grasp regions during the reverse diffusion process. In terms of grasp failures, we observe a major portion of them in large objects whose graspable regions are far from the center of mass. When lifted in simulation, the reaction torque snaps the object out of the gripper fingers and therefore is reported as a failure.
@@
Fig. 2(b) shows that our models also register low EMD for most test objects, demonstrating good coverage of grasp modes in the ground truth data. The EMD performance of GraspVAE and GraspLDM models are comparable or marginally better than the state-of-the-art grasp diffusion models from [9]. However, the results show that while GraspLDM models effectively learn the grasp distribution, they have a slightly worse EMD compared to the GraspVAE models. We attribute this to two factors. First, reverse diffusion moves poor grasp poses (e.g. colliding or free-space grasps) towards a smaller number of high-density regions that tend to provide a higher success rate. Second, we follow the computation of SE(3) EMD from [9] by using cosine distance for rotation and metric Euclidean distance for translation. As a result, the EMD metric is more sensitive to the similarity in rotation despite translation having a potentially larger effect on the success of grasp execution.
@@
Figure 5: Task-conditional de-noising in the latent space for a region-semantic class label (top, body, or bottom)
@@ IV-B Conditional generation
At test time, we generate 50 grasps for each mode (top, bottom, and body) of an object. The input object point cloud is augmented with random rotations. To check whether the generated grasp complies with the input label, we reverse the rotation transformation and check the location of the gripper’s root, as earlier. We report the precision between the labels of generated grasps and the labels supplied as conditioning. In this setting, the Task-GraspLDM model provides a mean precision of 0.703 averaged over all objects. We observe that the reverse diffusion process seamlessly moves a latent from a prior distribution to the desired task-conditional distribution, as visualized in Fig. 5. The tests show that GraspLDM allows effective injection of task conditioning post-hoc. Currently, the models are limited by the cases where there are no ground-truth grasps for a given label. For instance, if there are no ’bottom’ grasp annotations for a laptop, the model generates grasps unconditionally all over the object. This can be addressed in future work along with more complex task contexts.
@@
Sampling
$N_{G}=100$
$N_{G}=1000$
Median
success rate
DDPM
$7.39\pm 0.06$
$11.80\pm 0.29$
$\mathbf{0.792}$
DDIM
$\mathbf{0.75\pm 0.02}$
$\mathbf{1.14\pm 0.01}$
$0.756$
VAE
$\mathbf{0.02\pm 0.01}$
$\mathbf{0.03\pm 0.01}$
$0.660$
@@ IV-C Reverse diffusion sampling
A notable disadvantage of using DDPM-based grasp generation is the time to execute large number reverse sampling steps in a sequential manner. In DDPM, a large $T$ is required to ensure that the Gaussian conditional distributions in Eq. 4 are a good approximation [22]. A Denoising Diffusion Implicit Model [23] assumes a non-Markovian forward process to speed up the sample generation and requires a lower number of sampling steps. Further, it can be used as a drop-in sampler to use with a score network trained on the DDPM objective in Eq. 6, without any re-training. To assess the performance and sampling speed trade-off, we take the GraspLDM-63C model and compare DDIM sampler with 100 steps with the naive DDPM sampler. Table I compares the execution time of the reverse diffusion loop in DDIM with the baseline 1000-step DDPM. Using DDIM sampling as a drop-in replacement, the sampling time drops to 0.75s for 100 grasps and 1.1s for 1000 grasps. This is 10x faster than DDPM with a small loss in success rate (3.6%). The experiment was conducted on a system with NVIDIA RTX 3080Ti GPU and Intel i7-12800H CPU. Given the rapid progress in creating fast reverse diffusion samplers, our pipeline provides the flexibility to trade-off speed against performance for such diffusion samplers. In contrast, SE3-Grasp-DiF [9] models cannot leverage new samplers without re-training.
@@ IV-D Single-view point clouds
For many real-world use cases, only single-view point clouds are available for grasp generation. Therefore, we also evaluate the GraspLDM framework’s ability to learn the distribution of grasps on noisy partial point clouds. For evaluation, we use the simulation environment as before, with the addition of a depth camera. The cameras are spawned randomly between 30cm and 1m from the object. We use the partial point cloud thus obtained to sample 25 grasps and execute them without filtering. For each object, we repeat these for 20 random camera poses. The success rate is reported in Fig. 6 for GraspVAE-P-1C, GraspLDM-P-1C, GraspVAE-P-63C, and GraspLDM-P-63C models, where ’P’ implies that the model was trained on partial point clouds. In both cases, GraspLDM models improve the performance of the base VAE model. We found that for GraspLDM-P-63C models, higher capacity is required to compress a meaningful latent representation for 63C partial point clouds. We increase the grasp latent ($\mathbf{z}_{h}$) size to 16.
Note that we do not use preferential camera viewpoints, grasp filtering, or a support surface in the background. As a result, this evaluation provides a conservative metric for 6-DoF grasp-generation performance. This can be considered as the approximate lower bound of the grasp generation performance for real-world use cases. A large number of failures result from grasps that collide with parts of the objects not visible in the partial point clouds. Depending on the viewpoint, this occlusion confuses the model to see incomplete surfaces as edges along which the object could be grasped. For larger objects where grasps are concentrated in a small region of the object, these regions may not be available in a partial point cloud.
These issues highlight the need for discriminating bad grasps from good grasps for real-world tests. This can be done using a classification network on top of GraspLDM (as in real world tests in section IV-E) or preferably by adding a classification layer in the decoder. The latter problem of classifying grasps within the decoder of the current architecture, such that it can also be used for task-specific diffusion models, is out of the scope of the current work.
@@ IV-E Real robot tests
To complement the large-scale evaluations conducted in simulation, we conduct thorough real-world tests to evaluate the performance of object-centric GraspLDM models. We assess the transfer on a hardware setup that consists of a 6-DoF UR-10e arm, a Robotiq-3F gripper, and an Intel Realsense D435 RGB-D camera. The gripper is used in a two-finger mode with the gripper sweep length restricted to 85mm to emulate the gripper used for generating the training data. The camera is mounted at the tool flange of the arm in an eye-in-hand configuration. We use a test set of 16 objects of diverse physical and geometric properties as shown in Fig. 7. Each grasping trial is conducted by placing a single object in a random pose on a table. We conducted five random pose trials per object to evaluate the success rate across a total of 80 grasp attempts without retries. The robot is set to an arbitrary fixed pose such that the camera boresight is approximately ${}^{\circ}$ to the table plane and directed approximately to the center of the table.
@@
Method Success Rate
GraspLDM + GraspClassifier (ours) $\mathbf{80}$%
@@
6-DoF Graspnet + Classifier[3] $37.5$%
@@
We use the GraspLDM-P-63C model introduced in section IV-D without any further fine-tuning. Since the model provides object-centric grasps, we use the segment-anything model [24] to segment the object in an RGB image. Subsequently, we mask the aligned depth image to provide a segmented object point cloud. This point cloud is provided to the model which generates 100 grasp candidates. Unlike simulation, real experiments require a grasp pose to be selected. For this purpose, we train a naive classifier (GraspClassifier) to provide a pseudo-probability of success for a given pair of a point cloud and a grasp pose. Subsequently, we rank and sort the grasps above a probability threshold ($p>0.5$). We attempt the grasps in the same order while checking for collision and kinematic failures. The first grasp with a valid plan is executed on the robot. During a grasp attempt, the robot first goes to a pre-grasp pose, which is offset from the grasp pose along the z-axis of the gripper. Then the gripper approaches the object along this axis and closes the fingers. Subsequently, the object is lifted and moved to a set pose above a bucket. A grasp is successful if all these steps are executed successfully. Together these steps test the validity of the grasp pose as well as the stability of the grasp in all axes.
For motion planning, we utilize an off-the-shelf RRTConnect planner. We use 6-DoF-Graspnet [3] as the real-world test baseline as it is a widely used grasp generation model. For a fair evaluation of grasp synthesis models, we only use the generator and classifier from [3] and do not include the iterative refinement stage, while keeping all other parts of the pipeline including segmentation the same as before. We compare this with GraspLDM-P-63C and GraspVAE-P-63C in Table II. Grasp diffusion [9] is only trained on full point cloud models and therefore not used for comparison here.
Table II demonstrates that GraspLDM provides a superior grasp success rate in the real world while being entirely trained with simulation data. GraspLDM-P-63C model provides around 4.75% higher success rate over its base VAE model- GraspVAE-P-63C which also uses the same classifier. On the other hand, both our models provide a significantly higher success rate than 6-Dof-Graspnet [3] without its iterative refinement stage. We notice that high amounts of failure of the baseline result from grasps that collide with the object as well as from a lack of grasp generation diversity. In the latter case, it suffers from not producing enough grasps in kinematically feasible regions of the object in certain relative poses. We also observe a higher rate of planning failures for the baseline until a successful grasp is attempted, which is connected to grasp diversity. The results validate our architecture and training pipeline for practical applications. It is important to highlight that our models are only trained on 63 categories of a total of 180 in the ACRONYM dataset because of compute constraints. Therefore, we believe that the performance can be improved further.
V Conclusions

### v2

@@ IV Results and Discussion
We conduct a thorough evaluation of our models in four segments. First, the 6-DoF generation capability of our models is evaluated on full point clouds sampled from object meshes in Section IV-A. We then investigate the benefits of using latent diffusion by evaluating class-conditioned generation in Section IV-B and the flexibility of using a drop-in reverse diffusion sampler for speed improvements in Sec IV-C. In Section IV-D, we evaluate learning on partial point clouds from a single RGB-D image. Finally, we validate the practical effectiveness of our models and sim-to-real transfer with two real-world robotic setups in Section IV-E.
@@
Figure 4: Multi-object grasping environments with Franka gripper in Isaac Gym for success rate evaluation.
Figure 5: Grasp generation performance and scaling on full object point clouds ($N=1024$). (a) The mean success rate ina simulation of 300 generated grasps poses per object. (b) SE(3) EMD between ground-truth grasp pose distribution and 100 sampled grasp poses (lower is better). SE3-DiF-1C and SE3-DiF-63C are the SE(3) Grasp Diffusion models from [23].
@@
We evaluate the performance and scaling of GraspLDM on two category sets- 1C and 63C, where C denotes the number of ShapeNetSem categories used from the ACRONYM dataset. 1C is composed of 110 Mugs in the train set and 50 held-out Mugs in the test set. The 63C set contains 1100 objects from 63 categories in the train set and 400 held-out objects in the test set. In the case of both full and partial point clouds, the point cloud is randomly rotated, noised, and regularized to 1024 points during training and testing before it is fed to the model.
@@
We primarily evaluate the success rate of the grasp poses generated by the models. In Section IV-AIV-D, we assess the success rate with a large-scale parallel simulation in Isaac Gym [32]. Simulation offers the best method of exhaustively evaluating 6-DoF grasp generation. Real-world testing of 6-DoF grasps is limited by factors like kinematic feasibility, external collisions, and the time required per grasp execution. Therefore, we first evaluate the generator models in the simulation and subsequently validate the performance transfer to the real world. In the simulation, we use a two-fingered gripper (Franka hand) without the robotic arm as shown in Fig. 4 and ignore gravity to eliminate the effect of external factors on the evaluations. The simulation executes three steps for each grasp pose to report a grasp’s success or failure. (1) The gripper is spawned at the given grasp pose relative to the object with the fingers open. (2) The fingers are closed at the grasp pose. (3) The gripper shakes the object in X and Y directions and then lifts the object 1m in the +Z direction. The grasp is successful if the object remains attached to the gripper. Note that the evaluation is conservative as any penetration of the gripper in the first step or adverse contact in the second stage will result in a grasp failure. We observe that for large objects, ground truth grasp poses in the ACRONYM dataset provide a very low or zero success rate in our simulation. Therefore, we filter out the objects in each set for which the ground truth grasps fail more than 75% of the time.
@@
While the success rate evaluates the quality of the generated grasps, a model with a high success rate may generate grasps only around a small region of the object, instead of covering all the graspable regions around the object. Consequently, we also use SE(3) Earth Mover’s Distance (EMD) metric [23] to evaluate how well the trained models learn the object-centric grasp pose distribution. The SE(3) EMD metric evaluates the empirical distance between two distributions of SE(3) poses. To compute this, we sample 100 grasps per object from a model and sample 100 random grasps from the ground truth datasets. Subsequently, we compute the minimum distances between the two sets using linear sum assignment and report the mean distance over 500 iterations.
@@
For success rate, we outline two observations in Fig. 5(a). First, the LDM models consistently improve upon the base VAE models, demonstrating that the sample quality improves by using a diffusion model in VAE’s latent space. Second, the success rate performance of GraspLDM models scales better to a larger object set (63C). Comparing the median of the success rate on the test set in Fig. 5(a), GraspLDM-1C improves the median success rate to 88% compared to 84% of GraspVAE-1C. On the larger 63C set, the GraspLDM-63C model improves the median success rate to 78% compared to 66% from GraspVAE-63C. GraspLDM models also improve the interquartile range in each case. For the comparison against the baseline, SE3-DiF-1C model reports a median success rate of 89% which is comparable to 88% the GraspLDM-1C model. On the larger 63C set, we observe that the GraspLDM-63C model registers 78% median success rate and distinctly higher interquartile range compared to SE3-DiF-63C. Therefore, GraspLDM models scale more favorably to larger object sets compared to the state-of-the-art diffusion models for this task. During the tests, we also observed that the models available from [23] do not hold out a test split and are trained on all objects. Therefore, the comparison to SE3-DiF models [23] is conservative, which emphasizes the performance of GraspLDM on withheld objects.
Figure 6: Visualization of latent space denoising in GraspLDM. Reverse diffusion de-noises these latents from $t=T$ to $t=0$, gradually moving the bad grasps towards regions of good grasps.
@@
Overall, the results validate our hypothesis that a diffusion model can bridge the prior gap in the latent space to provide higher-quality grasp pose samples. Fig. 6 visualizes this effect where the latent denoising moves the latents corresponding to bad grasps towards good grasp regions during the reverse diffusion process. In terms of grasp failures, we observe a major portion of them in large objects whose graspable regions are far from the center of mass as shown in Fig. 11(a). When lifted in simulation, the reaction torque snaps the object out of the gripper fingers and therefore is reported as a failure.
@@
Fig. 5(b) shows that our models also register low EMD for most test objects, demonstrating good coverage of grasp modes in the ground truth data. The EMD performance of GraspVAE and GraspLDM models are comparable or marginally better than the SE3-DiF models. However, the results show that while GraspLDM models effectively learn the grasp distribution, they have a slightly worse EMD compared to the GraspVAE models. We attribute this to two factors. First, reverse diffusion moves poor grasp poses (e.g. colliding or free-space grasps) towards a smaller number of high-density regions that tend to provide a higher success rate. Second, we follow the computation of SE(3) EMD from [23] by using cosine distance for rotation and metric Euclidean distance for translation. As a result, the EMD metric is more sensitive to the similarity in rotation despite translation having a potentially larger effect on the success of grasp execution.
@@
Figure 7: Task-conditional denoising in the latent space for a region-semantic class label (top, body, or bottom)
@@ IV-B Conditional generation
Sampling $N_{G}=100$ $N_{G}=1000$ Median
(s) (s) success rate
DDPM $7.39\pm 0.06$ $11.80\pm 0.29$ $\mathbf{0.792}$
DDIM $\mathbf{0.75\pm 0.02}$ $\mathbf{1.14\pm 0.01}$ $0.756$
VAE $\mathbf{0.02\pm 0.01}$ $\mathbf{0.03\pm 0.01}$ $0.660$
At test time, we generate 50 grasps for each mode (top, bottom, and body) of an object. The input object point cloud is augmented with random rotations. To check whether the generated grasp complies with the input label, we reverse the rotation transformation and check the location of the gripper’s root, as earlier. We report the precision between the labels of generated grasps and the labels supplied as conditioning. In this setting, the Task-GraspLDM model provides a mean precision of 0.703 averaged over all objects. We observe that the reverse diffusion process seamlessly moves a latent from a prior distribution to the desired task-conditional distribution, as visualized in Fig. 7. These tests show that GraspLDM allows effective injection of task conditioning post-hoc. However, the models are limited by the cases where there are no ground-truth grasps for a given label. For instance, there are no ’bottom’ grasp annotations for a laptop. Consequently, the model generates grasps unconditionally all over the object as shown in Fig. 11(b) disregarding the label.
@@ IV-C Reverse diffusion sampling
A notable disadvantage of using DDPM-based grasp generation is the time to execute large number reverse sampling steps sequentially. In DDPM, a large $T$ is required to ensure that the Gaussian conditional distributions in Eq. 4 are a good approximation [18]. A Denoising Diffusion Implicit Model [33] assumes a non-Markovian forward process to speed up the sample generation and requires lower sampling steps. Further, it can be used as a drop-in sampler for a score network trained on the DDPM objective in Eq. 6, without any re-training. To assess the performance and sampling speed trade-off, we take the GraspLDM-63C model and compare the 100-step DDIM sampler against the naive 1000-step DDPM sampler. Table I compares the execution time of the reverse diffusion loop in DDIM with the baseline 1000-step DDPM. Using DDIM sampling as a drop-in replacement, the sampling time drops to 0.75s for 100 grasps and 1.1s for 1000 grasps. This is 10x faster than DDPM with a small loss in success rate (3.6%). The experiment was conducted on a system with NVIDIA RTX 3080Ti GPU and Intel i7-12800H CPU. Given the rapid progress in creating fast reverse diffusion samplers, our pipeline provides the flexibility to trade off speed against performance for such diffusion samplers. In contrast, SE3-Grasp-DiF [23] models cannot leverage new samplers without re-training.
@@ IV-D Single-view point clouds
For many real-world use cases, only single-view point clouds are available for grasp generation. Therefore, we also evaluate the GraspLDM framework’s ability to learn the distribution of grasps on noisy partial point clouds. For evaluation, we use the simulation environment as before, with the addition of a depth camera to simulate real-world conditions where only partial object information is available. The cameras are spawned randomly between 30cm and 1m from the object. We feed the partial point cloud obtained from this single RGB-D view to the models. We generate 25 grasps per point cloud and repeat the generation for 20 random camera poses per object. Subsequently, we test the 500 grasps per object in the simulation environment. The success rate over the two object sets is reported in Fig. 8 for GraspVAE-P-1C, GraspLDM-P-1C, GraspVAE-P-63C, and GraspLDM-P-63C models, where ’P’ implies that the model was trained on partial point clouds. In both cases, GraspLDM models improve the performance of the base VAE model. We found that for GraspLDM-P-63C models, higher capacity is required to compress a meaningful latent representation of partial point clouds from the 63 categories. This is expected as the model learning from partial point cloud needs to additionally reason about the camera pose and shape completion aspects to make sense of the grasps and learn a structured latent space. We therefore increase the grasp latent ($\mathbf{z}_{h}$) size to 16 for GraspLDM-P-63C.
Note that we do not use preferential camera viewpoints, grasp filtering, or a support surface in the background. As a result, this evaluation provides a conservative metric for 6-DoF grasp-generation performance. This can be considered as the approximate lower bound of the grasp generation performance for real-world use cases. A large number of failures result from grasps that collide with parts of the objects not visible in the partial point clouds. Depending on the viewpoint, this occlusion confuses the model to consider incomplete surfaces as edges along which the object could be grasped, as shown in Fig. 11(c). For larger objects where grasps are concentrated in a small region of the object, another issue affecting the performance is the visibility of these regions in the point cloud.
@@ IV-E Real robot tests
To complement the large-scale evaluations conducted in simulation, we conduct thorough real-world tests to evaluate the performance of object-centric GraspLDM models. We compare our performance against two established baselines widely used in the community. First, as a direct comparison with a generative model architecture, we use 6-DoF-Graspnet [13]. Second, we compare the effectiveness of our models against the state-of-the-art Contact-Graspnet [26], which uses a feed-forward PointNet style architecture. We do not use SE3-DiF diffusion models [23] for real-world tests as the training and testing in the original work are conducted with full object point clouds. A fair and reliable comparison with these models was not possible as we observed significant performance degradation when training SE3-DiF models on single-view point clouds from 63 categories with the default hyperparameters.
Figure 9: Examples of object placement on support during real-world testing. Such object placements allow grasps from more approach directions to be feasible while also demanding higher grasp pose accuracy for success.
Figure 10: Real-world validation of GraspLDM models with objects of diverse visual, geometric, and surface properties across two different robotic setups.
We assess the sim-to-real transfer on two hardware setups. The first setup uses a wall-mounted 6-DoF UR-10e arm, a Robotiq-3F gripper, and an Intel Realsense D435 RGB-D camera. The second setup uses a table-mounted 7-DoF Franka Research 3 arm with Franka hand. In the former case, the 3-finger gripper is operated in a two-finger ”pinch” mode with the gripper sweep length restricted to 85mm to emulate the gripper used for generating the training data. In both cases, the camera is mounted at the tool flange of the arm in an eye-in-hand configuration. Due to limitations of access to hardware, we conducted the comparison with 6-DoF Graspnet on the first setup, while the comparison with Contact-Graspnet was conducted later on the second setup. In each case, we use a test set of 16 randomly selected objects of diverse physical and geometric properties as shown in Fig. 10. Each grasping trial is conducted by placing a single object in a random pose on a table. To ensure that the trials allow execution of grasps from more approach directions than just top-down, we lift the objects on a small support as shown in Fig.  9. This also makes grasping success more difficult compared to table-top placement where in-plane sliding and normal contact force supports imprecise grasp attempts. At the start of the trial, the robot observes the object from an arbitrarily fixed pose such that the camera boresight is between 30 and 60 to the table plane and the object is between 0.4m and 0.8m from the camera. We conduct five random pose trials per object to evaluate the success rate across a total of 80 grasp attempts without retries.
@@
Setup Method Success
Rate
Setup 1: UR10e + Robotiq-3F GraspLDM+GraspClassifier (ours) $\mathbf{80}$%
@@
6-DoF-Graspnet+Classifier [13] $37.5$%
Setup 2: Franka FR3 + Franka hand GraspLDM+GraspClassifier (ours) $\mathbf{78.75}$%
GraspVAE+GraspClassifier (ours) $67.5$%
Contact-Graspnet [26] $76.25$%
@@
We use the GraspLDM-P-63C model introduced in Section IV-D without any further fine-tuning. Since the models provide object-centric grasps, we use the segment-anything model [34] to segment the object in an RGB image. Subsequently, we mask the aligned depth image to provide a segmented object point cloud. This point cloud is provided to the model which generates 100 grasp candidates. Unlike simulation, real experiments require a grasp pose to be selected. For this purpose, we train a naive classifier (GraspClassifier) to provide a pseudo-probability of success for a given pair of a point cloud and a grasp pose. Subsequently, we rank and sort the grasps above a probability threshold ($p>0.5$). We attempt the grasps in the same order while checking for collision and kinematic failures. The first grasp with a valid plan is executed on the robot. For motion planning, we utilize an off-the-shelf RRTConnect planner and plan cartesian trajectories to the grasp pose. During a grasp attempt, the robot first goes to a pre-grasp pose, which is offset from the grasp pose along the z-axis of the gripper. Then the gripper approaches the object along this axis and closes the fingers. Subsequently, the object is lifted and moved to a set pose above a bucket. A grasp is successful if all these steps are executed while keeping the object attached to the gripper. Together these steps test the validity of the grasp pose and the stability of the grasp in all axes. For a fair evaluation of grasp synthesis models, we only use the generator and classifier from 6-DoF-Graspnet and do not include the iterative refinement stage. Similarly, we do not perform iterative refinement on top of Contact-Graspnet predictions. All other parts of the pipeline like segmentation remain the same in every case. We compare this with GraspLDM-P-63C and GraspVAE-P-63C in Table II.
Table II demonstrates that GraspLDM provides a superior grasp success rate in the real world while being entirely trained with simulation data. Compared to the generative model baseline [13], GraspLDM demonstrates a significantly higher success rate. On the other hand, GraspLDM shows a comparable success rate with that of Contact-Graspnet [26], which is a feed-forward predictive model. Despite the success rate being marginally higher for GraspLDM, the difference concerns 2/80 grasp attempts. Due to several factors influencing real-world tests of 6-DoF grasps, we consider the two models to have comparable success rates. We also observe that the GraspLDM-P-63C model provides around 4.75% and 11.25% higher success rate over its base VAE model in the two test setups while using the same classifier. We believe that the relative variation in the performance of GraspVAE between the two test setups is mainly related to the non-trivial effects of viewpoint and noise on generation. We observe that for favorable viewpoints both GraspVAE and GraspLDM have similar output, while in some other cases, GraspLDM performs noticeably better than GraspVAE. This effect is more pronounced in the second setup than in the first.
We notice that high amounts of failure for the 6-DoF-Graspnet models result from the grasps colliding with the object and a lack of grasp generation diversity. In the latter case, it does not produce enough grasps in kinematically feasible regions of the object in a given pose. We also observe a higher rate of planning failures for this baseline until a successful grasp is attempted, which is connected to grasp diversity. On the other hand, Contact-Graspnet outperforms and produces superior grasps for objects closer to the categories represented in the ACRONYM dataset. This is especially the case for objects with trivial edges/surfaces like boxes. While for the objects with curvature and those with shapes unlike ACRONYM categories, it underperforms compared to GraspLDM. Overall, we demonstrate that GraspLDM models outperform existing generative models and provide comparable performance against larger feed-forward models.
Figure 11: GraspLDM Limitations: (a) Failure on large objects from adverse torque. (b) Failure of region-semantic conditional generation for ”bottom” grasps. (c) GraspLDM-P-63C generates colliding (orange) and free-space grasps (red).
V Limitations and Future Work
GraspLDM demonstrates that latent diffusion provides a powerful mechanism to improve generative performance, flexibility to downstream tasks, and sample quality for real-world grasping. However, this work is not without limitations. The primary limitation comes from not being able to bootstrap a notion of grasp quality. This currently necessitates a separate classifier in practical applications. In the real-world tests, we train a naive GraspClassifier, which is over-confident and prefers deeper grasp penetration around the object. An external classifier like this also hinders the framework’s flexibility to enable plug-and-play task-specific diffusion models which need to reason about the grasps that are stable and suitable for the task at the same time. Therefore, some notion of grasp quality should be systematically bootstrapped inside the GraspLDM architecture in the future.
Furthermore, our models are competitive despite being trained on only 63 out of 180 categories in the ACRONYM dataset due to compute constraints. We believe that the performance can be improved further by incorporating the entire dataset in the training. Finally, by training the latent diffusion model inside a frozen latent space of an unconditional GraspVAE, our framework paves the way for a flexible plug-and-play approach to task-specific grasping that inherits from a general grasping model. However, the analysis needs to be extended to more complex task conditioning inputs like points, poses, and embeddings.
VI Conclusions